OpenStudy (laddiusmaximus):

(sinx/1+cosx)+(1+cosx/sinx)=2cscx

myininaya (myininaya):

try to combine the fractions

OpenStudy (laddiusmaximus):

I was doing that. but It keeps telling me its wrong

OpenStudy (laddiusmaximus):

which one should be the common denominator?

OpenStudy (laddiusmaximus):

the sinx?

myininaya (myininaya):

is that suppose to be sinx/(1+cosx)+(1+cosx)/sinx

myininaya (myininaya):

to find a common denominator multiply the first fraction by sinx/sinx and the second fraction by (1+cosx)/(1+cosx)

OpenStudy (anonymous):

I think the common denominator would be (1+cosx)(sinx)

myininaya (myininaya):

so now you have [(sin)^2+1+2cosx+(cosx)^2]/[sinx(1+cosx)]=[1+1+2cosx]/[sinx(1+cosx)]=2(1+cosx)/[sinx(1+cosx)]=2/sinx=2cscx

OpenStudy (laddiusmaximus):

wait, what?

OpenStudy (laddiusmaximus):

im lost